Three equal charges 2×10^(-6) C each are held fixed at the three corners of an equilateral triangle of side 5cm. Find the Coulomb force experienced by one of the charges due to the rest two.

Grade:

2 Answers

Arun
5 years ago
The resultant force on charge Q because of two equal forces because of other two Q charges planed on equilateral triangle will be given as:-
Fnet = 2Fcos(ϴ/2) ---------------(1)
F=kQ1Q2/r2  --------(2)
in this question Q1=Q2 =Q3 =Q
Given
Q=2 *10-6 c
K=9*10 N-m2/c2
r= .05 m
ϴ =60 degree
By using equation 1 and 2 and given quantities
Fnet = 2(9*10*(2 *10-6 )2/(.05)cos30
=24.94 N
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem below.
 
The resultant force on charge Q because of two equal forces because of other two Q charges planed on equilateral triangle will be given as:-
Fnet = 2Fcos(ϴ/2) ---------------(1)
F = kQ1Q2/r2  --------(2)
in this question Q1 = Q2 = Q3 = Q
Given
Q = 2 *10-6 C
K = 9*10 N-m2/C2
r = .05 m
ϴ = 60o
By using equation 1 and 2 and given quantities
Fnet = 2(9*10*(2 *10-6 )2/(.05)cos30o
=24.94 N
 
Hope it helps.
Thanks and regards,
Kushagra